OpenStudy (anonymous):

Suppose a and b vary inversely, and b = 8 when a = 6. Write a function that models the variation and find b when a = 30.

hartnn (hartnn):

for inverse variation , the product of variables = constant. so here ab = k (a constant) so, can you find 'k' given that b=8,a=6 from ab=k ?

OpenStudy (anonymous):

Alright How would i start my problem ?

hartnn (hartnn):

i gave you start only. For, inverse variation. ab = k put a=6,b=8, then k=.... ?

OpenStudy (anonymous):

Oh okay. Hold on let me work it out

OpenStudy (anonymous):

do i do it like this (6)(8)=k ?

hartnn (hartnn):

yes, correct! so, k=... ?

OpenStudy (anonymous):

48

hartnn (hartnn):

correct :) so, your function is ab =48 . for 2nd part , just put a=30 in that function and find b.....

OpenStudy (anonymous):

when you mean put a for 30 isnt 6 already a unless i replace it

hartnn (hartnn):

yes, 8 and 6 are now useless. you have ab = 48 put a= 30 , what you get ?

OpenStudy (anonymous):

(30)(8) = 240

hartnn (hartnn):

8 is useless, keep b as b only so, 30 b=48 got this ? to get b now, divide both sides by 30..

OpenStudy (anonymous):

got that. Did what both sides by 30 ?

hartnn (hartnn):

divide

OpenStudy (anonymous):

30/48 ?

hartnn (hartnn):

30 b/30 = 48/30

OpenStudy (anonymous):

48/30 = 1.6

hartnn (hartnn):

so, b= 8/5 =1.6

OpenStudy (anonymous):

how you get b= 8/5 i thought it was 1/6

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

48 = 6*8 30 =6*5 48/30 = 6*8 /6*5 = 8/5

OpenStudy (anonymous):

Oh okay cool. Is this all for part 2 ? And is this all for my answer as well part 1 & 2

hartnn (hartnn):

yes, thats it... ab = 48 b= 8/5

OpenStudy (anonymous):

Thank you so much : )

hartnn (hartnn):

welcome ^_^